# MdEval: Massively Multilingual Code Debugging

### v1

@@ Abstract
Code large language models (LLMs) have made significant progress in code debugging by directly generating the correct code based on the buggy code snippet. Programming benchmarks, typically consisting of buggy code snippet and their associated test cases, are used to assess the debugging capabilities of LLMs. However, many existing benchmarks primarily focus on Python and are often limited in terms of language diversity (e.g., DebugBench and DebugEval). To advance the field of multilingual debugging with LLMs, we propose the first massively multilingual debugging benchmark, which includes 3.6K test samples of 18 programming languages and covers the automated program repair (APR) task, the code review (CR) task, and the bug identification (BI) task. Further, we introduce the debugging instruction corpora MdEval-Instruct by injecting bugs into the correct multilingual queries and solutions (xDebugGen). Further, a multilingual debugger xDebugCoder trained on MdEval-Instruct as a strong baseline specifically to handle the bugs of a wide range of programming languages (e.g. “Missing Mut” in language Rust and “Misused Macro Definition” in language C). Our extensive experiments on MdEval reveal a notable performance gap between open-source models and closed-source LLMs (e.g., GPT and Claude series), highlighting huge room for improvement in multilingual code debugging scenarios [^1] https://mdeval.github.io/.
@@ 1 Introduction
Large language models (LLMs) (OpenAI, 2023; Touvron et al., 2023a; Yang et al., 2024a) designed for code, such as CodeLlama (Rozière et al., 2023), DeepSeekCoder (Guo et al., 2024a), and QwenCoder (Hui et al., 2024a), are highly effective at code understanding and generation. These capabilities make them particularly useful for debugging, where deep comprehension of code structure and logic is essential. Automated program repair (APR) (Wen et al., 2024) aims to automatically fix software bugs without human involvement, significantly reducing time and costs in development processes.
LLMs have recently shown considerable potential in this area. For instance, CodeX (Chen et al., 2021) and GPT-4 series (OpenAI, 2023) outperforming previous conventional methods have demonstrated promising results on bug benchmarks like QuixBugs (Lin et al., 2017). The recent work DebugBench (Tian et al., 2024) creates a debugging benchmark including Python, Java, and CPP for LLM evaluation. However, for diverse programming languages in Figure 1, the multilingual debugging scenario poses more language-specific challenges for APR. Multilingual issues (e.g. “Misused Macro Definition” in programming language C, “Missing mut” in Rust, and “Unused Variable” in Go) highlight the complexities and diversities of locating and fixing bugs in the multilingual debugging scenario. Therefore, there is an urgent need to build a truly massively multilingual debugging code benchmark with a wide variety of generic and language-specific bug types.
@@
To further characterize the debugging performance of LLMs across different programming languages, we introduce MdEval, a framework for data construction, evaluation benchmark, and a multilingual debugging baseline xDebugCoder, to advance code debugging development. First, we propose MdEval, the first massively multilingual evaluation benchmark for code debugging covering 18 programming languages and 3.6K human-curated samples to assess the capabilities of LLMs across a wide range of languages. Further, we create MdEval-Instruct, a multilingual debugging instruction corpus in 18 languages to help the LLM fix the bug given the buggy code snippet. Specifically, we propose xDebugGen to create the buggy and correct code pair for debugging instruction tuning. The bugs are injected into the queries and solutions with our designed three strategies (1) Injecting bugs into query. (2) Injecting bugs into solution. (3) Injecting bugs with the round-trip code translation. Leveraging MdEval-Instruct, we develop xDebugCoder as a strong baseline, assessing the transferability of LLMs in multilingual debugging tasks.
The contributions are summarized as follows: (1) We propose MdEval, a comprehensive multilingual code debugging benchmark consisting of 3.6K samples spanning three tasks: automated program repair (APR), code review (CR), and bug identification (BI). This benchmark covers 18 languages and includes both generic and language-specific bug types. (2) We introduce the massively multilingual code debugging instruction corpora MdEval-Instruct created by the xDebugGen. By injecting bugs into the multilingual correct query or response, we can create the pairs of buggy code and the correct code for instruction. (3) We systematically evaluate the multilingual code debugging capabilities of 15+ models on our created MdEval and create a leaderboard to evaluate them on 18 programming languages dynamically. Notably, extensive experiments suggest that comprehensive multilingual multitask evaluation can realistically measure the gap between open-source (e.g. DeepSeekCoder and Qwen-Coder) and closed-source models (e.g. GPT-4o and Claude series).
@@ 2 Multilingual Code Debugging Evaluation: MdEval
Table 1: MdEval dataset statistics.
@@
Problems $3,513$
Automated Program Repair $1,171$
Code Review $1,171$
Bug Identification $1,171$
Total Test Cases $6,838$
@@
- Easy/Medium/Hard $1,692$/$1,209$/$612$
@@
      - maximum length $5913$ tokens
@@
      - avg length $239.4$ tokens
2.1 Data Overview
As illustrated in Table 1, the MdEval consists of 3.6K problems. Following Yang et al. (2024c), we design 3 multilingual debugging-related tasks: Automated Program Repair, Code Review, and Bug Identification. Each task contains about 1.2K questions, with more than 60 problems in each language. Each problem in MdEval includes question, example test cases, buggy code, correct code, and unit tests.
We calculated the token lengths of the question and buggy code using the CodeLlama tokenizer (Rozière et al., 2023). The average question length is 83 words, highlighting their detailed descriptive nature. The average buggy code length is 239 tokens, indicating the complexity of the code. In addition, the total number of unit tests for the dataset is 6,838, to ensure the accuracy of the bug-fix judgment. Figure 3 plots error types distribution. We strive to cover all error types in each language. Due to the inherent differences among languages, we ensure a balanced distribution of difficulty levels, leading to variations in the distribution of error types across languages.
(a) Generic Error Types
(b) Language-Specific Error Types
In Table 2, We compare MdEval with other code debugging benchmarks. Our benchmark provides a valuable enhancement to existing ones, significantly expanding the variety of programming languages and introducing language-specific error types, along with a greater number of questions and diverse bug-fixing tasks.
@@ 2.1 Data Overview
MdEval (Ours) 18 3 1,692/1,209/612 39 Human Annotation
@@ 2.2 MdEval Construction & Quality Control
We initially recruited 13 computer science graduates as multilingual debugging annotators, all proficient in their respective programming languages. After completing a comprehensive training course on annotation methods, the annotators were tasked with defining problems, providing corresponding solutions, and buggy code. Annotators adhered to the following principles: (1) Write a clear problem question and design test cases to ensure that bugs can be effectively identified; (2) Categorize bugs into multiple difficulty levels (easy/medium/hard) based on the complexity of fixing these code.
Figure 4 illustrates the overall process of dataset construction. We began by collecting code snippets from GitHub, which were then extracted and filtered following StarCoder (Li et al., 2023). Prior to the annotation phase, we summarized generic error types and language-specific error types. The three task definitions and corresponding annotation methods were explained in detail. The annotators proceeded to annotate the code according to the identified error types and specified annotation methods. To ensure annotation quality, they evaluated the annotated code based on four criteria: problem difficulty, ambiguity, error type, and solvable. Furthermore, after completing their annotations, each annotator exchanged data with another annotator for cross-refining, aiming to minimize subjective bias and errors. Any discrepancies between annotators were resolved through consensus or with input from senior annotators. Finally, we engaged three volunteers to assess the accuracy of the benchmark (targeting > 90%) and correct errors.
Figure 3: Error types distribution in 18 programming Languages from the MdEval.
@@
Figure 5: Examples of multilingual automated program repair (APR), code review (CR), and code identification (CI).
@@ 2.3 Instruction Corpora for Code Debugging
To create the instruction corpora, we need to create the pair of the correct code snippet and the buggy code. First, we select the proper code snippet from 18 languages and prompt the code LLM to generate a new question $q^{L_{k}}$ of programming language $L_{k}$. Then, we use the LLM to generate the correct code $c^{L_{k}}$ and filter the low-quality response with an LLM filter and the generated test cases. Therefore, we can regard the $(q^{L_{k}},c^{L_{k}})$ as the correct sample by ensuring the correctness of $c^{L_{k}}$ as much as possible. We propose xDebugGen comprised of the following three strategies to create the code debugging instruction corpora MdEval-Instruct to obtain the fine-tuned LLM xDebugCoder.
@@
We can prompt the LLM to modify the original question to another similar question with minor differences, where the similar question $q^{\prime L_{k}}$ is further used to generate the answer $\mathcal{M}_{w}(q^{\prime L_{k}})$ by a weak LLM $\mathcal{M}_{w}$ with small size (e.g. Qwen2.5-1.5B). Since there exist differences between the original question $q^{L_{k}}$ and the modified question $q^{\prime L_{k}}$, $(\mathcal{M}_{w}(q^{\prime L_{k}}),c^{L_{k}})$ can be fed into the LLM as source input and target prediction.
@@ 2.4 Evaluation Task
The automated program repair task forces the LLM to fix the bug in the given code snippet and then generates the correct code. Given the programming language $L_{k}\in\{L_{i}\}_{i=1}^{K}$ ($K=18$ is the number of programming languages), we provide the question $q^{L_{k}}$, the corresponding buggy code $b^{L_{k}}$, and the examples test cases $e^{L_{k}}$ for inputs. We can organize the different input settings for evaluation:
@@
Code Review (CR).
For the code review task, we present two versions of code to LLMs: the correct code $b^{L_{k}}$ and the buggy code $c^{L_{k}}$ with only a few minor differences between them. The correct code and buggy code are listed in a random order to feed into LLM for distinguishing the buggy code.
@@
In this task, LLMs are required to classify the type of error present in a given buggy program $c^{L_{k}}$ with one error. The LLMs must choose the correct error category from 39 bug types (including generic bug types and language-specific bug types).
@@ 2.5 Evaluation Metrics
In the automated program repair task, we evaluate models by executing the generated code against a set of unit tests and assessing performance using the Pass@1 metric (pass rate for just one-time generation). Greedy Pass@1 indicates whether a result produced by the LLM successfully passes all corresponding unit tests.
Code Review & Bug Identification.
In the code review and bug identification tasks, we evaluate model performance using accuracy, as both require the model to select from a set of provided options.
@@ 3.1 Experiment Setup
We evaluate 16 popular models, both closed-source and open-source (sizes ranging from 1.3B to 236B parameters). For general models, we evaluate GPTs (OpenAI, 2023) (GPT4-o, GPT4-o-mini), Claude-3.5 (Anthropic, 2023). For code models, we test Qwen2.5-Coder (Hui et al., 2024b), DeepSeekCoder (DS-Coder) (Guo et al., 2024a), CodeLlama (Rozière et al., 2023), and Codegemma (Gemma Team, 2024). Furthermore, we fine-tune the Qwen2.5-Coder-7B to provide a baseline model xDebugCoder for reference. For closed-source models, the responses are generated by the official API. For the open-source models, we perform inference on all models using the vLLM (Kwon et al., 2023) framework. All models adopt a greedy decoding strategy during inference, the temperature is set to 0, and the maximum generation length is 2048.
xDebugCoder Training Corpora.
xDebugCoder Optimization.
Our model, xDebugCoder, based on Qwen2.5-Coder-7B, is trained with full parameters for 3 epochs using a cosine scheduler, starting at a learning rate of $5\times 10^{-5}$ with 3% of total training steps for warmup. We utilize AdamW (Loshchilov and Hutter, 2017) as the optimizer; the batch size is set to 1024, with a maximum sequence length of 2048. All experiments are performed with 8 NVIDIA A800-80GB GPUs.
@@
Model Size Avgall C C# CLISP CPP F# Go JS Java
Closed-Source Models & 200B+ Models
[…]
DS-Coder-V2-Instruct 236B 72.7 60.9 87.1 70.0 63.6 80.0 44.6 61.4 78.3
1B+ Models
DS-Coder-Instrcut 1.3B 33.5 43.8 21.4 16.7 37.9 13.8 22.9 45.7 31.7
Yi-Coder-Chat 1.5B 33.0 35.9 38.6 8.3 39.4 9.2 26.5 45.7 36.7
Qwen2.5-Coder-Instruct 1.5B 36.2 29.7 2.9 13.3 33.3 16.9 10.8 58.6 35.0
6B+ Models
DS-Coder-Instruct 6.7B 52.9 59.4 7.1 53.3 63.6 23.1 37.3 67.1 60.0
CodeLlama-Instruct 7B 31.9 32.8 2.9 23.3 33.3 12.3 15.7 62.9 20
CodeGemma-Instruct 7B 48.9 45.3 2.9 3.3 36.4 20 41 71.4 66.7
CodeQwen1.5-Chat 7B 50.3 37.5 31.4 41.7 43.9 41.5 33.7 55.7 58.3
Qwen2.5-Coder-Instruct 7B 63.4 59.4 57.1 60.0 62.1 29.2 44.6 60.0 85.0
Yi-Coder-Chat 9B 48.3 54.7 7.1 16.7 47.0 16.9 4.8 68.6 80.0
DS-Coder-V2-Lite-Instruct 16B 59.5 60.9 12.9 40.0 57.6 32.3 32.5 65.7 76.7
DS-Coder-Instruct 33B 55.6 70.3 10 56.7 59.1 27.7 14.5 70 76.7
xDebugGen (Our Method) 7B 46.8 53.1 2.9 36.7 54.5 35.4 3.6 54.3 48.3
Method Json Julia PHP Pascal Python R Ruby Rust Scala Swift
Closed-Source Models & 200B+ Models
GPT-4o-240806 20.3 82.1 87.5 72.3 90.2 86.7 84.1 80.0 92.2 78.3
GPT-4o-mini-240718- 39.0 85.1 85.9 64.6 88.5 80.0 84.1 71.4 76.6 70.0
Claude-3.5-sonnet-240620 44.1 82.1 78.1 58.5 90.2 85.0 79.4 91.4 92.2 75.0
Claude-3.5-sonnet-241022 40.7 82.1 85.9 60.0 93.4 86.7 82.5 85.7 89.1 78.3
DS-Coder-V2-Instruct 44.1 83.6 71.9 67.7 85.2 75.0 81.0 85.7 87.5 86.7
1B+ Models
DS-Coder-Instrcut 8.5 46.3 31.2 40.0 49.2 20.0 57.1 32.9 31.2 53.3
Yi-Coder-Chat 16.9 40.3 28.1 24.6 54.1 20.0 52.4 42.9 34.4 36.7
Qwen2.5-Coder-Instruct 28.8 61.2 43.8 46.2 57.4 26.7 57.1 61.4 23.4 50.0
6B+ Models
DS-Coder-Instruct 23.7 67.2 57.8 58.5 78.7 48.3 82.5 71.4 35.9 63.3
CodeLlama-Instruct 11.9 43.3 34.4 38.5 39.3 21.7 63.5 58.6 23.4 36.7
CodeGemma-Instruct 28.9 68.7 48.4 50.8 73.8 46.7 74.6 75.7 64.1 61.7
CodeQwen1.5-Chat 23.7 64.2 53.1 52.3 68.9 45.0 71.4 77.1 37.5 71.7
Qwen2.5-Coder-Instruct 35.6 74.6 75.0 66.2 78.7 65.0 69.8 77.1 76.6 71.7
Yi-Coder-Chat 25.4 65.7 54.7 52.3 67.2 36.7 74.6 68.6 70.3 70.0
DS-Coder-V2-Lite-Instruct 30.5 70.1 68.8 64.6 73.8 71.7 77.8 82.9 79.7 83.3
DS-Coder-Instruct 28.9 70.1 54.7 50.8 73.8 60 85.7 67.1 76.6 63.3
xDebugGen (Our Method) 23.7 61.2 64.1 46.2 63.9 36.7 65.1 74.3 62.5 68.3
Table 4: Accuracy of different models for Code Review tasks on MdEval. The underlined numbers are the best scores for each language. “Avgall” represents the average accuracy of all code languages.
Model Size Avgall C C# CLISP CPP F# Go JS Java
@@
1B+ Models
DS-Coder-Instrcut 1.3B 54.5 68.8 57.1 51.7 49.2 62.1 45.8 58.6 53.3
Yi-Coder-Chat 1.5B 56.2 50.0 47.1 50.0 50.8 39.4 55.4 81.4 56.7
Qwen2.5-Coder-Instruct 1.5B 51.4 53.1 51.4 48.3 49.2 51.5 53.0 51.4 55.0

### v2

@@ Abstract
Code large language models (LLMs) have made significant progress in code debugging by directly generating the correct code based on the buggy code snippet. Programming benchmarks, typically consisting of buggy code snippets and their associated test cases, are used to assess the debugging capabilities of LLMs. However, many existing benchmarks primarily focus on Python and are often limited in terms of language diversity (e.g., DebugBench and DebugEval). To advance the field of multilingual debugging with LLMs, we propose the first massively multilingual debugging benchmark, which includes 3.9K test samples of 20 programming languages and covers the automated program repair (APR) task, the bug localization(BL) task, and the bug identification (BI) task. In addition, we introduce the debugging instruction corpora MdEval-Instruct by injecting bugs into the correct multilingual queries and solutions (xDebugGen). Further, a multilingual debugger xDebugCoder trained on MdEval-Instruct as a strong baseline specifically to handle bugs of a wide range of programming languages (e.g. “Missing Mut” in language Rust and “Misused Macro Definition” in language C). Our extensive experiments on MdEval reveal a notable performance gap between open-source models and closed-source LLMs (e.g., GPT and Claude series), highlighting huge room for improvement in multilingual code debugging scenarios [^1] https://mdeval-code.github.io/.
@@ 1 Introduction
Large language models (LLMs) (OpenAI, 2023; Touvron et al., 2023a; Yang et al., 2024a) designed for code, such as CodeLlama (Rozière et al., 2023), DeepSeekCoder (Guo et al., 2024a), and QwenCoder (Hui et al., 2024), are highly effective in code understanding and generation. These capabilities make them particularly useful for debugging, where deep comprehension of code structure and logic is essential. Automated program repair (APR) (Wen et al., 2024) aims to automatically fix bugs without human involvement, significantly reducing time and costs in development processes.
LLMs have recently shown considerable potential in this area. For instance, CodeX (Chen et al., 2021) and GPT-4 series (OpenAI, 2023) outperforming previous conventional methods have demonstrated promising results on bug benchmarks such as QuixBugs (Lin et al., 2017). The recent work DebugBench (Tian et al., 2024) creates a debugging benchmark including Python, Java, and CPP for LLM evaluation. However, for the diverse programming languages in Figure 1, the multilingual debugging scenario poses more language-specific challenges for APR. Multilingual issues (e.g. “Misused Macro Definition” in programming language C, “Missing mut” in Rust, and “Unused Variable” in Go) highlight the complexities and diversities of locating and fixing bugs in the multilingual debugging scenario. Therefore, there is an urgent need to build a truly massively multilingual debugging code benchmark with a wide variety of generic and language-specific bug types.
@@
To further characterize the debugging performance of LLMs across different programming languages, we introduce MdEval, a framework for data construction, evaluation benchmark, and a multilingual debugging baseline xDebugCoder, to advance the development of code debugging. First, we propose MdEval, the first massively multilingual evaluation benchmark for code debugging covering 20 programming languages and 3.9K samples to assess the capabilities of LLMs across a wide range of languages. Further, we create MdEval-Instruct, a multilingual debugging instruction corpus in 20 languages to help the LLM fix the bug given the buggy code snippet. Besides, we propose xDebugGen to create the buggy and correct code pair for debugging instruction tuning. The bugs are injected into the queries and solutions with our designed three strategies (1) Injecting bugs into query. (2) Injecting bugs into solution. (3) Injecting bugs with the round-trip code translation. Leveraging MdEval-Instruct, we develop xDebugCoder as a strong baseline, assessing the transferability of LLMs in multilingual debugging tasks.
The contributions are summarized as follows: (1) We propose MdEval, a comprehensive multilingual code debugging benchmark consisting of 3.9K samples spanning three tasks: automated program repair (APR), code localization (BL), and bug identification (BI). This benchmark covers 20 languages and includes both generic and language-specific bug types. (2) We introduce the massively multilingual code debugging instruction corpora MdEval-Instruct created by xDebugGen. By injecting bugs into the correct multilingual query or response, we can create pairs of buggy code and the correct code for instruction tuning. (3) We systematically evaluate the multilingual code debugging capabilities of 40 models on our created MdEval and create a leaderboard to evaluate them on 20 programming languages dynamically. Notably, extensive experiments suggest that comprehensive multilingual multitask evaluation can realistically measure the gap between open-source (e.g. DeepSeekCoder and Qwen-Coder) and closed-source models (e.g. Claude series).
@@ 2 Multilingual Code Debugging Evaluation: MdEval
2.1 Data Overview
In Table 1, the MdEval consists of 3.9K problems. Following Yang et al. (2024c), we design 3 multilingual debugging-related tasks: Automated Program Repair, Bug Localization, and Bug Identification. Each task contains about 1.3K questions, with more than 60 problems in each language. Each problem in MdEval includes question, example test cases, buggy code, correct code, and unit tests.
@@ 2.1 Data Overview
Problems $3,897$
Automated Program Repair $1,299$
Bug Localization $1,299$
Bug Identification $1,299$
Total Test Cases $7,133$
@@
- Easy/Medium/Hard $1,146$/$1,407$/$1362$
@@
      - maximum length $19,265$ tokens
@@
      - avg length $320.6$ tokens
Table 1: MdEval dataset statistics.
We calculate the length of the question and buggy code using the CodeLlama tokenizer (Rozière et al., 2023). The average question length is 83 words, highlighting their detailed descriptive nature. The average buggy code length is 239 tokens, indicating the complexity of the code. In addition, the total number of unit tests for the dataset is 6,838, to ensure the accuracy of the bug-fix judgment.
In Table 2, we compare MdEval with other code debugging benchmarks. Our benchmark provides a valuable enhancement to existing ones, significantly expanding the variety of programming languages and introducing language-specific error types, along with a greater number of questions and diverse bug-fixing tasks. The error types in MdEval are shown in Figure 2. Figure 3 plots error types distribution. We strive to cover all error types in each language. Due to the inherent differences among languages, we ensure a balanced distribution of difficulty levels, leading to variations in the distribution of error types across languages.
@@
MdEval (Ours) 20 3 1,692/1,209/612 47 Human Annotation
(a) Generic Error Types
(b) Language-Specific Error Types
@@ 2.2 MdEval Construction & Quality Control
We initially recruite 13 computer science graduates as multilingual debugging annotators, all proficient in their respective programming languages. After completing a comprehensive training course on annotation methods, the annotators are tasked with defining problems, providing corresponding solutions, and buggy code. Annotators adhere to the following principles: (1) Write a clear problem question and design test cases to ensure that bugs can be effectively identified; (2) Categorize bugs into multiple difficulty levels (easy/medium/hard) based on the complexity of fixing these code.
Figure 4 illustrates the overall process of dataset construction. We begin by collecting code snippets from GitHub, which are then extracted and filtered following StarCoder (Li et al., 2023). Prior to the annotation phase, we summarize generic error types and language-specific error types. The three task definitions and corresponding annotation methods are explained in detail. The annotators proceede to annotate the code according to the identified error types and specified annotation methods. To ensure annotation quality, they evaluate the annotated code based on four criteria: problem difficulty, ambiguity, error type, and solvability. Furthermore, after completing their annotations, each annotator exchanges data with another annotator for cross-refining, aiming to minimize subjective bias and errors. Any discrepancies between annotators are resolved through consensus or with input from senior annotators. Finally, we engage three volunteers to assess the accuracy of the benchmark (targeting > 90%) and correct errors.
Figure 3: Error types distribution in 20 programming Languages from the MdEval.
@@
Figure 5: Examples of multilingual automated program repair (APR), bug localization (BL), and code identification (CI).
@@ 2.3 Instruction Corpora for Code Debugging
To create the instruction corpora, we need to create the pair of the correct code snippet and the buggy code. First, we select the proper code snippet from 20 languages and prompt the code LLM to generate a new question $q^{L_{k}}$ of programming language $L_{k}$. Then, we use the LLM to generate the correct code $c^{L_{k}}$ and filter the low-quality response with an LLM filter and the generated test cases. Therefore, we can regard the $(q^{L_{k}},c^{L_{k}})$ as the correct sample by ensuring the correctness of $c^{L_{k}}$ as much as possible. We propose xDebugGen comprised of the following three strategies to create the code debugging instruction corpora MdEval-Instruct to obtain the fine-tuned LLM xDebugCoder.
@@
We can prompt the LLM to modify the original question to another similar question with minor differences, where the similar question $q^{\prime L_{k}}$ is used to generate the answer $\mathcal{M}_{w}(q^{\prime L_{k}})$ by a weak LLM $\mathcal{M}_{w}$ with small size (e.g. Qwen2.5-1.5B). Since there exist differences between the original question $q^{L_{k}}$ and the modified question $q^{\prime L_{k}}$, $(\mathcal{M}_{w}(q^{\prime L_{k}}),c^{L_{k}})$ can be fed into the LLM as source input and target prediction.
@@ 2.4 Evaluation Task
The automated program repair task forces the LLM to fix the bug in the given code snippet and then generates the correct code. Given the programming language $L_{k}\in\{L_{i}\}_{i=1}^{K}$ ($K=20$ is the number of programming languages), we provide the question $q^{L_{k}}$, the corresponding buggy code $b^{L_{k}}$, and the examples test cases $e^{L_{k}}$ for inputs. We can organize the different input settings for evaluation:
@@
Bug Localization (BL).
The Bug Localization (BL) task aims to identify the specific line(s) of code within a given buggy program $c^{L_{K}}$ that contains the error. For each test instance in the BL task, a buggy code $c^{L_{K}}$ is provided, from which four code snippets, $S_{A}$, $S_{B}$, $S_{C}$, $S_{D}$, are extracted. The LLMs are then tasked with identifying the golden snippet $S_{G}$, which contains the error.
@@
In this task, LLMs are required to classify the type of error present in a given buggy program $c^{L_{k}}$ with one error. The LLMs must choose the correct error category from 47 bug types (including generic bug types and language-specific bug types).
@@ 2.5 Evaluation Metrics
In the automated program repair task, we evaluate models by executing the generated code against a set of unit tests and assessing performance using the Pass@1 metric (pass rate for just one-time generation). Greedy Pass@1 indicates whether a result produced by the LLM successfully passes corresponding unit tests.
Bug Localization & Bug Identification.
In the bug localization and bug identification tasks, we evaluate model performance using accuracy, as both require the model to select from a set of provided options.
@@ 3.1 Experiment Setup
We evaluate 40 popular models, including GPTs (OpenAI, 2023), Claude-3.5 (Anthropic, 2023), and code-specific models like Qwen2.5-Coder (Hui et al., 2024), DeepSeekCoder (Guo et al., 2024a), CodeLlama (Rozière et al., 2023), and Codegemma (Gemma Team, 2024). Additionally, we fine-tune Qwen2.5-Coder-7B as our baseline xDebugCoder.
xDebugCoder Training Setup
The training data for xDebugCoder comprises our debugging dataset MdEval-Instruct and the Magicoder-Instruct code generation dataset (Wei et al., 2023), ensuring fundamental instruction-following capabilities for code-related tasks. xDebugCoder, built on Qwen2.5-Coder-7B, is trained for 3 epochs using a cosine scheduler with an initial learning rate of $5\times 10^{-5}$ with a 3% warmup ratio. We employ AdamW (Loshchilov and Hutter, 2017) as the optimizer, with a batch size of 1024 and a maximum sequence length of 2048. (Details can be found in the Appendix).
@@
Model Size Avgall C C# CLISP CPP F# Go HTML JS Java Json Julia MD PHP Pascal Python R Ruby Rust Scala Swift
@@
[…]
0.5B+ Models
Qwen2.5-Instruct 0.5B 20.6 28.6 10.4 8.3 14.5 9.2 1.2 13.8 45.3 28.3 10.2 26.9 5.0 17.4 13.8 39.3 6.7 58.7 24.3 18.8 31.7
DS-Coder-Instruct 1.3B 33.6 28.6 42.2 13.3 43.9 24.6 38.6 5.2 44.0 48.3 18.6 47.8 1.7 33.3 27.7 44.3 16.7 61.9 41.4 34.4 45.0
Qwen2.5-Instruct 1.5B 35.5 24.3 32.8 15.0 27.5 23.1 18.1 8.6 60.0 50.0 28.8 55.2 8.3 34.8 30.8 62.3 20.0 69.8 67.1 32.8 35.0
OpenCoder-Instruct 1.5B 34.8 15.7 13.4 20.0 26.1 26.2 15.7 12.1 57.3 58.3 15.3 55.2 8.3 36.2 47.7 54.1 31.7 68.3 52.9 51.6 28.3
Yi-Coder-Chat 1.5B 32.4 37.1 34.4 3.3 30.3 7.7 28.9 8.6 45.3 53.3 15.3 55.2 1.7 34.8 41.5 52.5 28.3 49.2 42.9 28.1 40.0
Qwen2.5-Coder-Instruct 1.5B 34.8 11.4 26.9 15.0 30.4 16.9 20.5 17.2 61.3 45.0 28.8 58.2 10.0 40.6 36.9 55.7 28.3 60.3 58.6 29.7 40.0
Qwen2.5-Instruct 3B 46.0 51.4 40.3 28.3 36.2 41.5 41.0 12.1 69.3 61.7 27.1 61.2 11.7 53.6 47.7 63.9 45.0 58.7 65.7 53.1 38.3